OpenStudy (anonymous):

what would be the slope intercept form of the line that is perpendicular to 7x+3y=18 and contains the point (6,8)

OpenStudy (anonymous):

slope intercept form is y = mx + b 7x + 3y = 18 => 3y = -7x + 18 7 => y = - --- x + 6 3 If a line is perpendicular to this, then its slope is negative reciprocal of the first line so the slope of the perpendicular line = 3/7 now m = 3/7, x= 6 and y =8 so 3 8 = ---- (6) + b 7 8 = 18/7 + b 8 - 18/7 = b 38/7 = b so the equation for perpendicular line will be 3 38 y = ---- x + ----- 7 7 7y = 3x + 38 or 3x - 7y = -38

OpenStudy (anonymous):

slope intercept form would be 3 38 y = ---- x + ----- 7 7